We will be back in UK waters shortly so underway sampling will start again. Everyone else is starting to pack up their equipment and looking forward to all the analysis of the frozen filters and samples when they get back to NOC.
Today has been a quiet day in that everyone has been bent over their laptops securing their data, and log sheets and writing their sections of the cruise reports.
Today was programmed to be the last on site, so there was lots to be packed in. The last Megacore was completed over night, and the PELAGRAs floats were recovered after 48 hours deployed.
Time out here is precious so the nights are busy too. Corinne deployed a new zooplankton net overnight; the last one got ripped as it was hauled back on board and the Mega Corer was deployed again.
So far we have had remarkably benevolent weather and yesterday I caught the sun a little as I was on deck filming most of the day, so today I started by putting sun screen on! It has been needed too as I have spent a good deal of today on deck filming.
After the actual recovery of the mooring buoy and then the sensor frame, the long job of recovering the 4km of rope began and didn't finish until late in the evening. It was not until 22:30 that three PELAGRA buoys could be deployed on a trial mission.